two planes started at the same time from the same airport and flew in opposite directions. one flew 60 miles per hour faster than the other. in 5 hours, they were 2800 miles apart. find the rate of each plane.
Answer by mananth(12270)

Plane I x mph
Plane II x+ 60
..
They are moving away from each other

So add up their speed.

combined speed = x+ 1 x + 60
( 2 x + 60 )
Time = 5 hours
Distance = 2800 miles
Distance = speed * time
( 2 x 60 )* 5 = 2800
10 x + 300 = 2800
10 x = 2800 + -300
10 x = 2500
/ 10
x= 250 mph Plane I
Plane II 250 + 60 = 310 mph